OpenStudy (anonymous):

please help Last year, Stacy purchased 81 shares of Stock A at $55 per share, 64 shares of Stock B at $57 per share, and a ten-year $3500 bond with a 6.16% coupon for $3100. Stacy sold both stocks today. Stock A is worth $61 per share and Stock B has a value of $63 per share. Assuming neither stock paid a dividend, which investment has the highest rate of return Stock A Stock B Bond Stock A and Stock B

OpenStudy (anonymous):

@mathstudent55 @NaCl

OpenStudy (mathstudent55):

Stock A: 81 shares bought last year Bought at $55 per share. Sold at $61 per share \( \dfrac{$61 - $55}{$55} \times 100 = 10.9\%\) Stock B: 64 shares bought last year Bought at $57 per share. Sold at $63 per share \( \dfrac{$63 - $57}{$57} \times 100 = 10.5\%\) Sorry, I don't know how bonds work.

OpenStudy (anonymous):

what do i do from here?

OpenStudy (mathstudent55):

Find out what the bond coupon percentage means. Does that mean the bond gives a 6.16% rate of return? If so, then the highest rate of return was stock A with 10.9% rate of return. Like I wrote above, I don't know how bonds work.